# Nonparametric Uniform Inference in

### v1

@@ Abstract
We develop methods for nonparametric uniform inference in cost-sensitive binary classification, a framework that encompasses maximum score estimation, predicting utility maximizing actions, and policy learning. These problems are well known for slow convergence rates and non-standard limiting behavior, even under point identified parametric frameworks. In nonparametric settings, they may further suffer from failures of identification. To address these challenges, we introduce a strictly convex surrogate loss that point-identifies a representative nonparametric policy function. We then estimate this surrogate policy to conduct inference on both the optimal classification policy and the optimal policy value. This approach enables Gaussian inference, substantially simplifying empirical implementation relative to working directly with the original classification problem. In particular, we establish root-$n$ asymptotic normality for the optimal policy value and derive a Gaussian approximation for the optimal classification policy at the standard nonparametric rate. Extensive simulation studies corroborate the theoretical findings. We apply our method to the National JTPA Study to conduct inference on the optimal treatment assignment policy and its associated welfare.
@@ 1 Introduction
If a researcher further wishes to allow for a nonparametric policy class to accommodate more flexible classifiation policies, the problem becomes even more severe. Parametric problems are still point-identified after suitable normalizations (e.g., Manski, 1975, 1985). By contrast, there is no obvious normalization that guarantees point identification when the policy function is nonparametric. Statistical inference under partially identified nuisance parameters poses significant challenges, and the development of valid and practical inference methods in this setting remains an open question.
This paper proposes a novel method of Gaussian inference for cost-sensitive binary classification problems with a nonparametric class of classification policy functions. The framework encompasses the aforementioned maximum score estimator, utility maximization under binary action, and treatment choice problems, among others. While this may seem daunting, if not infeasible, to those familiar with the literature, we attempt to develop a formal framework that establishes the validity of the proposed Gaussian inference procedure and demonstrate its practical ease.
@@
Once the representing nonparametric optimizer is “identified,” we can proceed with econometric analyses for point-identified functional parameters, even though the underlying model admits infinitely many parameters in the identified set. We estimate this nonparametric representative policy function using a sieve approach. Our procedure goes beyond standard sieve M-estimation by accommodating infinite-dimensional nuisance parameters, such as the propensity score and conditional mean functions arising in the augmented inverse propensity score weighting estimator for example. In this setting, we achieve the standard nonparametric rate of convergence. Moreover, we establish a strong Gaussian approximation for the representative policy by applying Yurinskii’s coupling inequality (Pollard, 2002, Theorem 10), which in turn enables Gaussian inference for the optimal classification policy.
With an estimator for the optimal representative policy, a plug-in yields an estimator of the functions of the optimal policy, such as the optimal policy values. We establish the root-$n$ convergence in distribution to the Gaussian limit for this plug-in estimator. This result is achieved through a combination of several techniques and notable properties of our problem to be outlined ahead, and it shares a similar spirit to Chen et al. (2003), who derive root-$n$ semiparametric results in related settings with non-smooth objective functions. In our setting, we show that a margin condition (see Kitagawa and Tetenov, 2018; Ponomarev and Semenova, 2024) plays a crucial role in ensuring that the population value function is sufficiently smooth at the optimal policy, thereby laying the foundation for Gaussian inference for the optimal policy values. As we accommodate infinite-dimensional nuisance parameters in the estimation of a finite-dimensional parameter, our approach bears resemblance to double machine learning (DML, Chernozhukov et al., 2018). However, the asymptotic analysis differs from DML because we additionally employ a nonparametrically estimated functional optimizer, i.e., the classification policy.
@@
Parametric Kitagawa and Tetenov (2018) & Athey and Wager (2021), etc. Rai (2018)
@@
As already noted, the (A) maximum score literature (e.g., Kim and Pollard, 1990; Horowitz, 1992) develops limit distributions for inference under parametric classification functions. In addition, Rai (2018) develops hypothesis tests for (B) empirical welfare maximization frameworks under finite VC dimensions. We contribute to this literature by allowing the VC dimension of the classes of classification policy functions to diverge to infinity, so as to accommodate nonparametric policy functions.
@@
In the context of treatment assignment, several plug-in approaches to inference exist, leveraging nonparametric nuisance parameters, such as the conditional average treatment effects, that are directly estimable. Bhattacharya and Dupas (2012) construct confidence intervals for the optimal welfare under budget constraints. Armstrong and Shen (2023) develop inference methods for decision boundaries. Park (2025) derives asymptotic normality by smoothing the indicator function à la Horowitz (1992). In contrast, our approach is aligned with empirical welfare maximization as a special case of cost-sensitive binary classification, and is therefore complementary to these plug-in methods. [^6] See Hirano and Porter (2020, Section 2.3) for details about the conceptual distinction between the plug-in approach and the empirical welfare maximization approach in statistical decisions. Furthermore, our general binary classification framework extends beyond treatment assignment settings to encompass maximum score and utility maximization frameworks.
@@
Finally, it is worth noting several related results from the statistics and biostatistics literature. Luedtke and Van Der Laan (2016) develop root-$n$ rate confidence intervals for the optimal welfare even when the parameter is not pathwise differentiable. Wu and Wang (2021) propose a smoothing-based procedure for model-free inference on the optimal treatment. Liang et al. (2022) employ surrogate convex losses to conduct inference on low-dimensional parameters. Wu et al. (2023) develop a uniform inference method by quantifying the unknown link function, thereby enabling flexible modeling of interactions between treatment and covariates. The present paper, which adopts a surrogate-based approach to uniform inference for nonparametric policies, complements specific aspects of each of these contributions.
@@ 2 Setup and Examples
Let $Z_{i}$ and $X_{i}$ denote random vectors observed by researchers or policymakers. Some variables may be common to both $Z_{i}$ and $X_{i}$. Let $\mathcal{G}$ denote a class of functions. With these notations, a broad class of econometric frameworks can be represented as a cost-sensitive binary classification problem:
@@
$L(g)=\mathbb{E}\!\left[\,\psi^{+}(Z_{i},\eta)\,\mathbf{1}\{g(X_{i})\geq 0\}+\psi^{-}(Z_{i},\eta)\,\mathbf{1}\{g(X_{i})<0\}\right].$ (2.1)
@@ 2.1 Example I: Maximum Score
where the error term $v_{i}$ satisfies $\text{Median}(v_{i}|X_{i})=0$ a.s. The maximum score estimator (Manski, 1975, 1985) for $g$ is defined as the sample analogue of
@@ 2.2 Example II: Expected Utility Maximization with Binary Actions
$\max_{a(\cdot)}\mathbb{E}\left[U(a(X),Y,X)\right]$
under a $\{-1,1\}$-valued binary outcome $Y$ with respect to a binary action $a(x)\in\{-1,1\}$ for each $x$. Imposing the condition that U(1,1,X)>U(1,1,X)U(1,1,X)>U(-1,1,X) and U(1,1,X)>U(1,1,X)U(-1,-1,X)>U(1,-1,X) to ensure correct classification yields higher utility, Elliott and Lieli (2013) show that the problem can be reformulated as $\max_{g\in\mathcal{G}}V(g)$, where
$V(g)=\mathbb{E}[b(X)[Y+1-2c(X)]\cdot\mathbf{1}\{g(X)\geq 0\}-b(X)[Y+1-2c(X)]\cdot\mathbf{1}\{g(X)<0\}],$
b(x)=U(1,1,x)U(1,1,x)+U(1,1,x)U(1,1,x)b(x)=U(1,1,x)-U(-1,1,x)+U(-1,-1,x)-U(1,-1,x), and c(x)=[U(1,1,x)U(1,1,x)]/b(x)c(x)=[U(-1,-1,x)-U(1,-1,x)]/b(x). The optimal action $a^{*}(x)$ can then be recovered from the optimal classifier $g^{\ast}$ as $a^{\ast}(x)=\mathrm{sign}(g^{*}(x))$ for each $x$. Like Example I (Section 2.1), the above maximization problem can be further reformulated as a special case of the general risk-minimization framework with the risk given in (2.1) $\blacktriangle$
@@ 3 Issues
The primary challenge arises from the discontinuous nature of the loss function. In practice, this discontinuity induces computational difficulties: the sample objective function is non-convex, making numerical optimization difficult, if not infeasible. Beyond these practical concerns, the discontinuity also has adverse theoretical implications, giving rise to non-standard limiting behavior and non-identification, as detailed in the following two subsections.
@@ 3.1 Non-Standard Limit under Identifying Normalizations
To simplify the discussion, suppose for the moment that $g$ belongs to a parametric class specified as $g(x)=x^{\prime}\gamma$, with appropriate normalizations. In this case, the problem $\min_{g\in\mathcal{G}}L(g)$ reduces to an $M$-estimation of $\gamma$ with a discontinuous objective function. In this setting, Kim and Pollard (1990) show that the resulting $M$-estimator converges at the cube-root-$n$ rate and admits a non-standard asymptotic distribution. Consequently, naive bootstrap procedures are invalid in this context (Abrevaya and Huang, 2005; Léger and MacGibbon, 2006).
@@ 4 Surrogate Identification
$L^{\ast}(g):=\mathbb{E}\left[\psi^{+}(Z_{i},\eta)\cdot\phi(-g(X_{i}))+\psi^{-}(Z_{i},\eta)\cdot\phi(g(X_{i}))\right].$ (4.1)
@@
Suppose that $\mathbb{E}[\psi^{+}(Z_{i},\eta)|X_{i}]-\mathbb{E}[\psi^{-}(Z_{i},\eta)|X_{i}]\neq 0$ almost surely and $\mathcal{G}$ includes all measurable functions. Let $\phi(\cdot)$ be differentiable at zero and $\partial\phi(0)<0$.
@@
Although we identify only one representative solution $g_{\ast}$, this suffices for conducting inference on the optimal classification policy. Indeed, $g_{\ast}$ is sign-equivalent to any solution of $\min_{g\in\mathcal{G}}L(g)$ up to a null set. Accordingly, the optimal classification rule defined by $x\mapsto\mathbf{1}\{x\in\mathcal{X}\mid g_{\ast}(x)\geq 0\}$ delivers the same policy decisions as the rule $x\mapsto\mathbf{1}\{x\in\mathcal{X}\mid g(x)\geq 0\}$ associated with any solution $g\in\arg\min_{g\in\mathcal{G}}L(g)$, possibly except on a set of measure zero. Thus, we may perform inference on the optimal classification policy and the corresponding optimal policy value (e.g., welfare) using the representative, point-identified surrogate policy function $g_{\ast}$, effectively treating it as the unique solution.
@@
then the function $g$ could take either positive or negative values on that event, thereby failing to point-identify the solution. It is worth noting that we allow $\mathbb{E}[\psi^{+}(Z_{i},\eta)\mid X_{i}]$ and $\mathbb{E}[\psi^{-}(Z_{i},\eta)\mid X_{i}]$ to be two distinct constants (excluding the uninteresting case where they coincide). This condition is essential not only for point identification but also for the inference on the optimal policy value; see, for example, Luedtke and Van Der Laan (2016, Theorem 1) for related discussions.
5 Estimation and Inference for the Surrogate Policy Function
Given the ‘identification’ result from Section 4, one can obtain a point estimate of the representing, surrogate policy function $g_{\ast}$. Although $g_{\ast}$ is only one among infinitely many solutions, it suffices for constructing the optimal classification rule $x\mapsto\mathbf{1}\{g_{\ast}(x)\geq 0\}$, since only the sign of $g_{\ast}$ matters and any other solution has an equivalent sign almost everywhere. Consequently, the optimal classification policy and the optimal policy value (e.g., welfare) can be estimated.
5.1 Estimation of the Surrogate Policy Function
One may adopt alternative nonparametric approaches to estimate the surrogate policy function $g_{\ast}$, but for the convenience of readers considering practical applications, we provide a concrete procedure.
@@ 5.1 Estimation of the Surrogate Policy Function
For a given fold $\ell\in[m]=\{1,\ldots,m\}$, obtain an estimator $\widehat{\eta}_{(\ell)}$ of the nuisance parameter $\eta_{0}$ using the complementary subsample $I_{(\ell)}^{c}$. Then, estimate the surrogate policy function $g_{\ast}$ on the held-out fold $I_{(\ell)}$. Specifically, the fold-wise estimator of $g_{\ast}$ is given by
@@
5.2 Large Sample Theory for Surrogate Policy Function Estimation

### v2

@@ Abstract
We develop methods for nonparametric uniform inference in cost-sensitive binary classification, a framework that encompasses maximum score estimation, predicting utility maximizing actions, and policy learning. These problems are well known for slow convergence rates and non-standard limiting behavior, even under point identified parametric frameworks. In nonparametric settings, they may further suffer from failures of identification. To address these challenges, we introduce a strictly convex surrogate loss that point-identifies a representative nonparametric policy function. We then estimate this representative policy function to conduct inference on both the optimal classification policy and the optimal policy value. This approach enables Gaussian inference, substantially simplifying empirical implementation relative to working directly with the original classification problem. In particular, we establish root-$n$ asymptotic normality for the optimal policy value and derive a Gaussian approximation for the optimal classification policy at the standard nonparametric rate. Extensive simulation studies corroborate the theoretical findings. We apply our method to the National JTPA Study to conduct inference on the optimal treatment assignment policy and its associated welfare.
@@ 1 Introduction
If a researcher further wishes to allow for a nonparametric policy class to accommodate more flexible classifiation policies, the problem becomes even more severe. Parametric problems are still point-identified after a suitable normalization (e.g., Manski, 1975, 1985). By contrast, there is no obvious normalization that guarantees point identification when the policy function is nonparametric. Statistical inference under partially identified parameters poses significant challenges, and the development of valid and practical inference methods in this setting remains an open question.
This paper proposes a novel method of Gaussian inference for cost-sensitive binary classification problems with a nonparametric class of classification policy functions. The framework encompasses the aforementioned maximum score estimation, utility maximization under binary action, and treatment choice problems, among others. While this may seem daunting, if not infeasible, to those familiar with the literature, we attempt to develop a formal framework that establishes the validity of the proposed Gaussian inference procedure and demonstrate its practical ease.
@@
Once the representing nonparametric optimizer is “identified,” we can proceed with econometric analyses for point-identified functional parameters, even though the underlying model admits infinitely many parameters in the identified set. We estimate this nonparametric representative policy function using a sieve approach. Our procedure goes beyond standard sieve M-estimation by accommodating infinite-dimensional nuisance parameters, such as the propensity score and conditional mean functions arising in the augmented inverse propensity score weighting estimator. In this setting, we achieve the standard nonparametric rate of convergence. Moreover, we establish a strong Gaussian approximation for the representative policy by applying Yurinskii’s coupling inequality (Pollard, 2002, Theorem 10), which in turn enables Gaussian inference for the optimal classification policy.
With our estimator for the optimal representative binary classification policy, we can estimate the optimal policy value by evaluating its (cross-fitted) empirical analog at our optimal policy estimator. We establish its root-$n$ consistency and limiting Gaussian distribution. This result is achieved through a combination of several techniques and notable properties of our problem to be outlined ahead, and it shares a similar spirit to Chen et al. (2003), who derive root-$n$ semiparametric results in related settings with non-smooth objective functions. In our setting, we show that a margin condition (see Kitagawa and Tetenov, 2018; Ponomarev and Semenova, 2024) plays a crucial role in ensuring that the population policy value function is sufficiently smooth at the optimal policy, thereby laying the foundation for Gaussian inference for the optimal policy values. As we accommodate infinite-dimensional nuisance parameters in the estimation of a finite-dimensional parameter, our approach bears resemblance to double machine learning (DML, Chernozhukov et al., 2018). However, the asymptotic analysis differs from DML because we additionally employ a nonparametrically estimated functional optimizer, i.e., the classification policy.
@@
Parametric Kitagawa and Tetenov (2018) & Athey and Wager (2021), etc. Rai (2018), etc.
@@
As already noted, in the maximum score literature (A), Kim and Pollard (1990) and Horowitz (1992) develop limit distributions of the orignal maximum score estimator and the smoothed maximum estimator, respectively. Chen and Lee (2018) construct a best subset maximum score prediction rule that is linear in covariates by maximizing a penalized objective function. Recently, Chen et al. (2025) reformulate the maximum score estimator by rectified linear unit (ReLU) functions to achieve computation advantage and convergence rate that is faster than cube-root-n. All these work consider a linear index of finite number of covariates. Under the empirical welfare maximization framework (B), Rai (2018) conduct inference on optimal policy functions and maximized welfare under finite VC dimension. Chernozhukov et al. (2025) introduce a framework for selecting policies that maximize a penalized empirical welfare function with the penalty being proportional to a consistent estimator of the risk of each candidate policy. They primarily focus on cases in which the set of policies has finite carnality, with brief discussion of extensions to infinity. We contribute to the both literature, (A) and (B), by allowing the VC dimension of the classes of classification policy functions to diverge to infinity, so as to accommodate nonparametric policy functions.
@@
In the context of treatment assignment, several plug-in approaches to inference exist, leveraging nonparametric nuisance parameters, such as the conditional average treatment effects, that are directly estimable (e.g., Bhattacharya and Dupas, 2012; Armstrong and Shen, 2023; Park, 2025; Chen et al., 2025). In contrast, our approach is aligned with empirical welfare maximization as a special case of cost-sensitive binary classification, and is therefore complementary to these plug-in methods. [^6] See Hirano and Porter (2020, Section 2.3) for details about the conceptual distinction between the plug-in approach and the empirical welfare maximization approach in statistical decisions. Our approach is also conceptually different from Xu (2025), who analyzes the optimal point decision from a continuum of alternatives in limit experiments. Furthermore, our general binary classification framework extends beyond treatment assignment settings to encompass maximum score and utility maximization frameworks.
@@
Finally, it is worth noting several related results from the statistics and biostatistics literature. Luedtke and Van Der Laan (2016) develop root-$n$ rate confidence intervals for the optimal welfare even when the parameter is not pathwise differentiable. Wu and Wang (2021) propose a smoothing-based procedure for model-free inference on the optimal treatment. Liang et al. (2022) employ surrogate convex losses to conduct inference on low-dimensional parameters. Wu et al. (2023) develop a uniform inference method by quantifying the unknown link function, thereby enabling flexible modeling of interactions between treatment and covariates. The present paper, which adopts a surrogate-loss approach to uniform inference for nonparametric policies, complements specific aspects of each of these contributions.
The paper is organized as follows. Section 2 introduces the setup and presents motivating examples. Section 3 discusses the key issues arising from the discontinuous nature of the loss function. Section 4 develops the “identification” of the representative classification policy function. Section 5 presents estimation and inference procedures for the representative classification policy function, and Section 6 establishes estimation and inference results for the optimal policy value. Section 7 examines the finite-sample performance of the proposed methods through simulation studies. Section 8 provides an empirical analysis that demonstrate the application of our method. Section 9 concludes.
@@ 2 Setup and Examples
Let $Z_{i}$ and $X_{i}$, $i=1,2,\cdots,n$, denote random vectors of unit $i$ observed by researchers or policymakers. Some variables may be common to both $Z_{i}$ and $X_{i}$. Let $\mathcal{G}$ denote a class of functions. With these notations, a broad class of econometric frameworks can be represented as a cost-sensitive binary classification problem:
@@
$L(g)=\mathbb{E}\!\left[\,\psi^{+}(Z_{i},\eta_{0})\,\mathbf{1}\{g(X_{i})\geq 0\}+\psi^{-}(Z_{i},\eta_{0})\,\mathbf{1}\{g(X_{i})<0\}\right].$ (2.1)
@@ 2.1 Example I: Maximum Score
where the error term $v_{i}$ satisfies $\text{Median}(v_{i}|X_{i})=0$ almost surely. The maximum score estimator (Manski, 1975, 1985) for $g$ is defined as the sample analogue of
@@ 2.2 Example II: Expected Utility Maximization with Binary Actions
$\max_{a(\cdot)}\mathbb{E}\left[U(a(X_{i}),Y_{i},X_{i})\right]$
under a $\{-1,1\}$-valued binary outcome $Y_{i}$ with respect to a binary action $a(x)\in\{-1,1\}$ for each $x$. Imposing the condition that U(1,1,x)>U(1,1,x)U(1,1,x)>U(-1,1,x) and U(1,1,x)>U(1,1,x)U(-1,-1,x)>U(1,-1,x) for all $x$ in the support of $X$ to ensure correct classification yields higher utility, Elliott and Lieli (2013) show that the problem can be reformulated as $\max_{g\in\mathcal{G}}V(g)$, where
$V(g)=\mathbb{E}[b(X_{i})[Y_{i}+1-2c(X_{i})]\cdot\mathbf{1}\{g(X_{i})\geq 0\}-b(X_{i})[Y_{i}+1-2c(X_{i})]\cdot\mathbf{1}\{g(X_{i})<0\}],$
b(x)=U(1,1,x)U(1,1,x)+U(1,1,x)U(1,1,x)b(x)=U(1,1,x)-U(-1,1,x)+U(-1,-1,x)-U(1,-1,x), and c(x)=[U(1,1,x)U(1,1,x)]/b(x)c(x)=[U(-1,-1,x)-U(1,-1,x)]/b(x). The optimal action $a^{*}(x)$ can then be recovered from the optimal classifier $g^{\ast}$ as $a^{\ast}(x)=\mathrm{sign}(g^{*}(x))$ for each $x$. Like Example I (Section 2.1), the above maximization problem can be further reformulated as a special case of the general risk-minimization framework with the risk given in (2.1) by taking $Z_{i}:=(Y_{i},X_{i}^{\prime})^{\prime}$, $\psi^{+}(z,\eta):=|b(x)(y+1-2c(x))|-b(x)(y+1-2c(x))$, and $\psi^{-}(z,\eta):=|b(x)(y+1-2c(x))|+b(x)(y+1-2c(x))$. $\blacktriangle$
@@ 3 Issues
The primary challenge arises from the discontinuous nature in the sample analog of the objective function. In practice, this discontinuity induces computational difficulties: the sample objective function is non-convex, making numerical optimization difficult, if not infeasible. Beyond these practical concerns, the discontinuity also has adverse theoretical implications, giving rise to non-standard limiting behavior and non-identification, as detailed in the following two subsections.
@@ 3.1 Non-Standard Limit under Identifying Normalizations
To simplify the discussion, suppose for the moment that $g$ belongs to a parametric class specified as $g(x)=x^{\prime}\gamma$, with appropriate normalizations. In this case, the problem $\min_{g\in\mathcal{G}}L(g)$ reduces to an $M$-estimation of $\gamma$ with a discontinuous sample objective function. In this setting, Kim and Pollard (1990) show that the resulting $M$-estimator converges at the cube-root-$n$ rate and admits a non-standard asymptotic distribution. Consequently, naive bootstrap procedures are invalid in this context (Abrevaya and Huang, 2005; Léger and MacGibbon, 2006).
@@ 4 Representative Identification
$L^{\ast}(g):=\mathbb{E}\left[\psi^{+}(Z_{i},\eta_{0})\cdot\phi(-g(X_{i}))+\psi^{-}(Z_{i},\eta_{0})\cdot\phi(g(X_{i}))\right].$ (4.1)
@@
Suppose that $\mathbb{E}[\psi^{+}(Z_{i},\eta_{0})|X_{i}]-\mathbb{E}[\psi^{-}(Z_{i},\eta_{0})|X_{i}]\neq 0$ almost surely and $\mathcal{G}$ includes all measurable functions. Let $\phi(\cdot)$ be differentiable at zero and $\partial\phi(0)<0$.
@@
Although we identify only one representative solution $g_{\ast}$, this suffices for conducting inference on the optimal classification policy. Indeed, $g_{\ast}$ is sign-equivalent to any solution of $\min_{g\in\mathcal{G}}L(g)$ up to a null set. Accordingly, the optimal classification rule defined by $x\mapsto\mathbf{1}\{x\in\mathcal{X}\mid g_{\ast}(x)\geq 0\}$ delivers the same policy decisions as the rule $x\mapsto\mathbf{1}\{x\in\mathcal{X}\mid g(x)\geq 0\}$ associated with any solution $g\in\arg\min_{g\in\mathcal{G}}L(g)$, possibly except on a set of measure zero. Thus, we may perform inference on the optimal classification policy and the corresponding optimal policy value (e.g., welfare) using the representative, point-identified representative policy function $g_{\ast}$, effectively treating it as the unique solution.
@@
The use of different loss functions leads to different representers $g_{*}$. However, as argued above, these representers yield the same optimal classification rule, since they agree in sign except on a null set. As such, the choice of loss function, and hence the resulting representer, may be viewed as analogous to selecting different parameter normalizations proposed by Manski (1975, 1985) in the context of the maximum score framework (Example I) introduced in Section 2.1.
@@
then the function $g$ could take either positive or negative values on that event, thereby failing to point-identify the solution. This condition is essential not only for point identification but also for the inference on the optimal policy value; see, for example, Luedtke and Van Der Laan (2016, Theorem 1) for related discussions.
5 Estimation and Inference for the Representative Policy Function
Given the ‘identification’ result from Section 4, one can obtain a point estimate of the representative classification policy function $g_{\ast}$. Although $g_{\ast}$ is only one among infinitely many solutions, it suffices for constructing the optimal classification rule $x\mapsto\mathbf{1}\{g_{\ast}(x)\geq 0\}$, since only the sign of $g_{\ast}$ matters and any other solution has an equivalent sign almost everywhere. Consequently, the optimal classification policy and the optimal policy value (e.g., welfare) can be estimated.
5.1 Estimation of the Representative Policy Function
@@
For a given fold $\ell\in[m]=\{1,\ldots,m\}$, obtain an estimator $\widehat{\eta}_{(\ell)}$ of the nuisance parameter $\eta_{0}$ using the complementary subsample $I_{(\ell)}^{c}$. Then, estimate the representative policy function $g_{\ast}$ on the held-out fold $I_{(\ell)}$. Specifically, the fold-wise estimator of $g_{\ast}$ is given by
@@
5.2 Large Sample Theory for the Representative Policy Function Estimation